The lengths of pregnancies of humans are normally distributed with a mean of 268 days and a standard deviation of 15 days. A baby is premature if it is born three weeks early. What percentage of babies are born prematurely?
Answer by khwang(438)

3 weeks = 21 days.
z = -21/15 = -7/5.
From the table of normal distribution , the prob. is
0.08

Kenny